A store mixes Kentucky bluegrass worth $11 per pound and che wings fescue worth$14 per pound. The mixture is to sell for $13 per pound. Find how much of each should be used to make a 420- pound mixture.
How many pounds of the Kentucky bluegrass should be in the mixture?
How many points of the che wings fescue should be in the mixture?

Answer by richwmiller(17219)

11*k+14f=13*420,
k+f=420
f = 280, k = 140